The following information is taken from the accounts of Latta Company. The entries in the T-accounts are summaries of the transactions that affected those accounts during the year.

    

Manufacturing Overhead Work in Process

   (a)              460,000

  (b)              390,000  

  Bal.            15,000

(c)              710,000  

260,000  

Bal.              70,000  

  

85,000  

      
          

(b)               390,000  

      
          
      

  Bal.               40,000  

      

  

Finished Goods Cost of Goods Sold

Bal.              50,000

  (d)              640,000

(d)              640,000   

  

(c)                710,000

                  

                   

  
  

Bal.               120,000

   

  

      

  

The overhead that had been applied to production during the year is distributed among the ending balances in the accounts as follows:

    

    
  Work in process, ending $ 19,500
  Finished goods, ending 58,500
  Cost of goods Sold 312,000
  
  Overhead applied $ 390,000
   

  

For example, of the $40,000 ending balance in work in process, $19,500 was overhead that had been applied during the year.

   

Required:
1. Identify the reasons for entries (a) through (d).

          

2.

Assume that the company closes any balance in the manufacturing overhead account directly to cost of goods sold. Prepare the necessary journal entry. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

             

3.

Assume instead that the company allocates any balance in the manufacturing overhead account to the other accounts in proportion to the overhead applied during the year that is in the ending balance in each account. Prepare the necessary journal entry. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Do not round intermediate calculations.)
Answer #1
Concepts and reason

Accounting: Accounting is the process of recording the transactions, classifying it in the specific manner and further the process of summarising and, analysing is done and thus the results are interpreted. It is the process of preserving the accounts.

Transaction: Transaction is an act of buying or selling goods or rendering any service that is reliably measured in terms of money.

Journal entry: Journal entry is the recording of transactions in a systematic manner as they occur. Thus it is a summary of all transactions which has debit and credit aspects recorded chronologically.

Fundamentals

Overheads: Overheads are the indirect expenses borne by the company. The types of overheads are Factory overheads, office and administration overheads and selling and distribution overheads.

Work in Process: Work in process is the partially completed units. It is in the stage before completion.

Manufacturing Overhead: Manufacturing overheads are the overheads related to the manufacturing of product. They are the indirect expenses during manufacturing.

Finished goods: Finished goods are the goods completed and being ready for sale. They are the stock and regarded as current assets of the company.

Cost of Goods Sold: Cost of Goods Sold is the cost that is directly incurred for the sale of goods. It includes cost such as travelling expenses during the time of sale, salesmen remuneration, etc. It is also known as cost of sales.

Rules for debit and credit

When asset increases debit it; asset decreases credit it.

When liabilities increases credit it; liabilities decreases debit it.

When stockholders’ equity increases credit it; stockholders’ equity decreases debit it.

Expenses and losses increases debit it; expenses and losses decreases credit it.

Incomes and gains increases credit it; incomes and gains decreases debit it

(1.a)

Entry (a) denotes the actual manufacturing overhead cost for the period and it is the balance brought down during the period.

(1.b)

The balance of $390,000 represents the actual manufacturing cost applied to production during the period. Thus it is credited in the manufacturing overheads account

(1.c)

The cost of completed goods is $710,000 which has been transferred to finished goods inventory from the work in process account.

(1.d)

In the cost of goods sold account it represents the finished goods inventory and it has been transferred to cost of goods sold account. It represents the cost of goods sold for the year.

(2)

Journal entry for applied manufacturing overhead

Date
Credit
Debit
$70,000
Account title and explanation
Cost of goods sold
Manufacturing overhead
(To close the under applied

(3)

Journal entry

Date
Credit
Account title and explanation
Work in process (19,500/390,000)*70,000
Finished goods (58,500/390,000)*70,000
Cost

Working notes:

AmountallocatedtoWorkinprocess=($19,500$390,000)×70,000=$3,500\begin{array}{c}\\{\rm{Amount allocated to Work in process = }}\left( {\frac{{\$ 19,500}}{{\$ 390,000}}} \right) \times 70,000\\\\{\rm{ = \$ 3,500}}\\\end{array}

Therefore amount allocated to Work in process is $3,500

AmountallocatedtoCostofgoodssold=($312,000$390,000)×$70,000=$56,000\begin{array}{c}\\{\rm{Amount allocated to Cost of goods sold = }}\left( {\frac{{\$ 312,000}}{{\$ 390,000}}} \right) \times \$ 70,000\\\\{\rm{ = \$ 56,000}}\\\end{array}

Therefore amount allocated to Cost of goods sold is $56,000

AmountallocatedtoFinishedgoods=($58,500$390,000)×$70,000=$10,500\begin{array}{c}\\{\rm{Amount allocated to Finished goods = }}\left( {\frac{{\$ 58,500}}{{\$ 390,000}}} \right) \times \$ 70,000\\\\{\rm{ = \$ 10,500}}\\\end{array}

Therefore amount allocated to finished goods is $10,500

Ans: Part 1.a

The actual manufacturing overhead during the period is $460,000

Part 1.b

It represents the manufacturing overhead cost applied to production during the period.

Part 1.c

It represents the cost of goods manufactured during the year

Part 1.d

It represents the cost of finished goods inventory which has been transferred to cost of goods sold.

Part 2

Date
Credit
Debit
$70,000
Account title and explanation
Cost of goods sold
Manufacturing overhead
(To close the under applied

Part 3

Date
Credit
Account title and explanation
Work in process
Finished goods
Cost of goods sold
Manufacturing overhead
(To close